OpenStudy (anonymous):

Simplify the quantity x plus 6 over 12, all over the quantity x minus 8 over 10.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@Coolsector

OpenStudy (studygurl14):

Like this? \(\Large\frac{x+\frac{6}{12}}{x-\frac{8}{10}}\) Or like this: \(\Large\frac{\frac{x+6}{12}}{\frac{x-8}{10}}\)

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

@uri

OpenStudy (studygurl14):

Okay. You want to eliminate the denominator and put everything in the numerator. To do that, you need to multiply the denominator by the its reciprocal so that it equals 1, and can be "eliminated". (Since any number with 1 as a denominator can be written without the denominator. Ex. \(\frac{8}{1}=8\)). However, to prevent changing the value of the fraction, we also have to multiply the numerator by the denominator's reciprocal. See here: \(\huge\frac{\frac{x+6}{12}}{\frac{x-8}{10}}\times\frac{\frac{10}{x-8}}{\frac{10}{x-8}}\rightarrow\frac{\frac{(x+6)10}{12(x-8)}}{\frac{x-8(10)}{10(x-8)}}\rightarrow\frac{\frac{(x+6)10}{12(x-8)}}{1}\rightarrow\frac{(x+6)10}{12(x-8)}\) Can you do the rest? @shira_lew

OpenStudy (studygurl14):

Oops, got cut off.

OpenStudy (studygurl14):

\(\Large\rightarrow\frac{(x+6)10}{12(x-8)}\)